Calculate the following and show all the calculations.
5\6 ÷ 1\12=?

5\8 × 1\2=?

2 1\3 × 1 3\4=?

2\5 ÷ 3=?
1
Expert's answer
2020-07-06T20:23:21-0400

Solution. Using the rules for performing arithmetic operations with fractions, we obtain


"\frac {5}{6} \div \frac {1}{12}=\frac {5}{6}×\frac{12}{1}=\frac{60}{6}=10""\frac {5}{8} × \frac {1}{2}=\frac {5}{16}"

"2\frac {1}{3} ×1 \frac {3}{4}=\frac {7}{3} × \frac {7}{4}=\frac{49}{12}=4 \frac{1}{12}"


"\frac {2}{5} \div 3=\frac {2}{5}×\frac{1}{3}=\frac{2}{15}"


Answer.

"\frac {5}{6} \div \frac {1}{12}=10"

"\frac {5}{8} × \frac {1}{2}=\frac {5}{16}"

"2\frac {1}{3} ×1 \frac {3}{4}=4 \frac{1}{12}"


"\frac {2}{5} \div 3=\frac{2}{15}"